OpenStudy (anonymous):

PLEASE HELP! See the three. What proportion of the first 1000 natural numbers have a 3 somewhere in them? For example, 135, 403, and 339 all contain a 3, whereas 402, 677, and 8 do not.

OpenStudy (anonymous):

@ganeshie8 hi do you think you could help me with this? i don't understand it at all!

OpenStudy (dangerousjesse):

How many numbers between 0 and 100 contain a 3?

OpenStudy (dangerousjesse):

It shouldn't take very long to count them.. :/

OpenStudy (dangerousjesse):

I'm just going to assume you know by now, the amount is x. \(x\times 10+89=\textsf{the answer}\) The 89 is for all 300's, excluding the ones already counted in \(\times 10\)

OpenStudy (anonymous):

but it says to 1000?

OpenStudy (anonymous):

@ali1029 still need help?